List of row vectors and column vectors of the matrix.




[2 -1 0 1



3 5 7 -1



1 4 2 7]






1
Expert's answer
2022-06-06T08:34:58-0400
"A=\begin{bmatrix}
 2 & -1 & 0 & 1 \\
 3 & 5 & 7 & -1 \\
 1 & 4 & 2 & 7 \\
\end{bmatrix}"

Row vectors


"\begin{bmatrix}
 2 & -1 & 0 & 1
 
\end{bmatrix},\begin{bmatrix}
 3 & 5 & 7 & -1
 
\end{bmatrix}, \begin{bmatrix}
 1 & 4 & 2 & 7
 
\end{bmatrix}"

Column vectors


"\begin{bmatrix}
 2 \\
 3 \\
 1 \\
\end{bmatrix}, \begin{bmatrix}
 -1 \\
 5 \\
 4 \\
\end{bmatrix}, \begin{bmatrix}
 0 \\
 7 \\
 2 \\
\end{bmatrix}, \begin{bmatrix}
 1 \\
 -1 \\
 7 \\
\end{bmatrix}"